OpenStudy (anonymous):

Where does the graph of y = 2x2 − x − 3 cross the x-axis?

OpenStudy (anonymous):

Crossing the x-Axis means \[f(x)=0\ \\ \therefore 2x^2-x-3 = 0 \] Which is the characteristic quadratic equation. \[ \frac{-b \pm \sqrt{D}}{2a} \] with \[D= b^2-4ac\]

OpenStudy (anonymous):

so what is b, a and c?

OpenStudy (anonymous):

The characteristic quadratic equation looks like this: \[ f(x)=ax^2+bx+c\] so in your case \[a=2, b=-1, c=-3\]

OpenStudy (anonymous):

alright thank you!